a plane can cost accomodate 325 passengers for an overseas flight. a first class seat cost $750 and a economy sear cost $525. the fuel for the flight cost $3000. if the revenue for ticket sales is $193 125 for a full plain, find the number of first class seats and economy seats.
Answer by edjones(7311)

Let f=number of first class passengers; 325-f=number of economy seats.
750f+525(325-f)=193125
750f+170625-525f=193125
225f=22500
f=100
325-100=225 economy seats.
.
Ed